### v1

# Entity-aware Cross-lingual Claim Detection for Automated Fact-checking

Abstract
Identifying claims requiring verification is a critical task in automated fact-checking, especially given the proliferation of misinformation on social media platforms. Despite significant progress in the task, there remain open challenges such as dealing with multilingual and multimodal data prevalent in online discourse. Addressing the multilingual challenge, recent efforts have focused on fine-tuning pre-trained multilingual language models. While these models can handle multiple languages, their ability to effectively transfer cross-lingual knowledge for detecting claims spreading on social media remains under-explored. In this paper, we introduce EX-Claim, an entity-aware cross-lingual claim detection model that generalizes well to handle claims written in any language. The model leverages entity information derived from named entity recognition and entity linking techniques to improve the language-level performance of both seen and unseen languages during training. Extensive experiments conducted on three datasets from different social media platforms demonstrate that our proposed model significantly outperforms the baselines, across 27 languages, and achieves the highest rate of knowledge transfer, even with limited training data.
Introduction
Automated fact-checking is an emerging research task dedicated to combating misinformation, particularly prevalent on social media, and comprises several key stages: claim detection, claim prioritization, evidence retrieval, and claim validation (Zeng, Abumansour, and Zubiaga 2021). While research in the task is progressing rapidly, identifying and validating claims related to global concerns requires a fact-checking pipeline capable of processing claims written in multiple languages. However, developing multilingual solutions for fact-checking research is not straightforward due to the availability of limited training data, especially for low-resource languages (Panchendrarajan and Zubiaga 2024). This necessitates advancements in multilingual fact-checking research to leverage limited training data for developing generalized solutions that can effectively transfer knowledge between languages.
This research focuses on the first component of the fact-checking pipeline: verifiable claim detection. A verifiable claim is defined as a statement expressing facts, excluding personal experience and private knowledge (Panchendrarajan and Zubiaga 2024). Existing works on verifiable claims detection are predominantly focused on monolingual solutions (Prabhakar, Mohtaj, and Möller 2020; Suri and Dudeja 2022; Henia et al. 2021; Hussein, Ghneim, and Joukhadar 2021), with limited attention given to addressing the multilingual nature of the problem. Especially, fine-tuning multilingual language models such as mBERT (Alam et al. 2021b; Uyangodage, Ranasinghe, and Hettiarachchi 2021; Panda and Levitan 2021) and XLM-R (Alam et al. 2021b; Hüsünbeyi, Deck, and Scheffler 2022; Savchev 2022; Alam et al. 2021a) is observed as a common solution. While these models can handle multiple languages, none of them adequately validate the cross-lingual knowledge-transferring capability of these models, particularly in identifying claims written in languages that were not seen during training. To address this challenge, we propose EX-Claim, a cross-lingual claim detection model that can effectively identify verifiable claims regardless of their language. Our model broadens the ability to perform across a diverse set of languages beyond those seen in training by incorporating entity-centric information for knowledge transfer. Our approach leverages the limited multilingual data available in claim detection research to develop a robust cross-lingual model that generalizes beyond the languages seen in the training phase.
We build our model on the assumption that factual claims circulating in social media often revolve around entities. In particular, entity types and their statuses such as their popularity may affect the verifiability of a claim. For instance, consider the following three statements.
  • S1: X visited Denmark to attend ICWSM
  • S2: Keir Starmer visited France to attend Euro Con
  • S3: Keir Starmer visited his home in the UK
Statement $S1$ can be deemed unverifiable from the perspective of a fact-checker, as it describes a personal experience of an unknown individual X. However, a similar context of attending an event expressed in $S2$ is verifiable as the popular entity Keir Starmer (the UK Prime Minister at the time of writing) attending an official event Euro Con becomes verifiable. However, changing the popular entities Keir Starmer or Euro Con with another popular person or event entity will not change the verifiability status of this claim. At the same time, changing the event with a private location such as home reduces the verifiability of the claim $S3$ even if a popular entity is involved. This motivates our intuition that entities, their types, their popularity, and the relationship between them affect the verifiability status of the claims.
Our proposed EX-Claim builds on this intuition by leveraging entity information to effectively identify verifiable claims in a cross-lingual environment. Our model incorporates entity information such as type and popularity derived from named entity recognition (NER) and entity linking (EL). Extensive experiments conducted using three datasets from different social media platforms comprising 27 languages show that EX-Claim significantly outperforms the baselines and reaches the highest knowledge transfer rate between languages seen and unseen in training. We make the following key contributions:
  • We propose EX-Claim, a cross-lingual claim detection model for identifying verifiable claims from social media which can effectively transfer knowledge across languages seen and unseen during training.
  • We enhance language-level performance by leveraging entity information in claims through NER and EL.
  • We conduct extensive experiments to evaluate EX-Claim across three datasets representing different social media platforms and encompassing 27 languages, including synthetic data created via machine translation.
The source code and the synthetic data are available online [^1] https://github.com/RubaP/Ex-Claim.
Related Work
A common approach to identifying multilingual claims shared on social media involves fine-tuning pre-trained multilingual language models such as mBERT (Alam et al. 2021b; Uyangodage, Ranasinghe, and Hettiarachchi 2021; Panda and Levitan 2021; Zengin, Kartal, and Kutlu 2021; Hasanain and Elsayed 2022) and XLM-R (Alam et al. 2021b, a). Most of these works (Alam et al. 2021b, a; Uyangodage, Ranasinghe, and Hettiarachchi 2021; Baris-Schlicht, Magnossao de Paula, and Rosso 2021; Du, Gollapalli, and Ng 2022) fine-tune the language model using combined training data of multiple languages and evaluate the capability of the model in identifying claims written in the same set of languages. Very few studies (Panda and Levitan 2021; Zengin, Kartal, and Kutlu 2021; Hasanain and Elsayed 2022) explore the model’s ability to transfer knowledge across language pairs, and these studies are limited to a narrow selection of languages.
Rather than focusing on developing cross-lingual methods, several existing studies have aimed at creating language-specific models to achieve optimal performance. For instance, Hüsünbeyi, Deck, and Scheffler (2022); Savchev (2022); Eyuboglu et al. (2023) fine-tuned models such as mBERT or XLM-R using language-specific training data. With the recent attention to large language models (LLM), Agrestia, Hashemianb, and Carmanc (2022) utilized GPT-3 to fine-tune the model in each language to develop a language-specific claim detection model. Interestingly, the results revealed that transformer-based architectures like BERT remain competitive with, and in some cases outperform the large language models (LLMs). Following these findings, further research (Sawiński et al. 2023; Lewoniewski et al. 2024; Li, Panchendrarajan, and Zubiaga 2024) explored leveraging LLMs for claim detection by modeling the task as a text generation problem. While these models demonstrate superior performance in the English language, their ability to handle multilingual claims for the claim detection task has yet to be explored.
Apart from the direct utilization of multilingual language models, various other techniques have been employed to address the challenges posed by limited training data and the noisy nature of social media data. This includes the extraction of platform-specific textual features (Alam et al. 2021b), data augmentation techniques such as machine translation (Zengin, Kartal, and Kutlu 2021; Savchev 2022) and sampling (Zengin, Kartal, and Kutlu 2021) and multitasking approaches (Baris-Schlicht, Magnossao de Paula, and Rosso 2021; Du, Gollapalli, and Ng 2022). While these techniques are explored in multilingual environments as well, none of them have been shown to improve cross-lingual knowledge-transferring capabilities, which our study performs comprehensively for the first time looking at 27 different languages through the exploration of entity-based knowledge transfer.
Figure 1: EX-Claim : Entity-aware Cross-lingual Claim Detection
Methodology
Given a statement written in any language, we classify it as verifiable or non-verifiable. Figure 1 illustrates the proposed solution, EX-Claim: entity-aware cross-lingual claim detection framework, whose components are discussed next.
Cross-lingual Word Embedding
We exploit various multilingual pre-trained language models to extract embeddings from their final layer. Especially, we experiment with XLM-R (Kalyan, Rajasekharan, and Sangeetha 2021), and mBERT (Devlin et al. 2019) to obtain the cross-lingual vector representation of words. XLM-R was trained on CommonCrawl data supporting 100 languages, and mBERT was trained on Wikipedia data containing 104 languages. Both models generate embeddings of size 768 for each tokenized word.
Named Entity Recognition (NER)
Identifying entities mentioned in the text is essential to extract further information such as its type and popularity. We use state-of-the-art NER models WikiNEuRal (Tedeschi et al. 2021) and MultiNERD (Tedeschi and Navigli 2022) for identifying coarse-grained and fine-grained entities respectively. We use the WikiNEuRal model and fine-tune the MultiNERD model with recommended parameters using their source code and annotated dataset. WikiNEuRal recognizes 4 types of coarse-grained named entities and MultiNERD recognizes 15 types of fine-grained named entities. Refer to Appendix A and A for the named entity types and statistics of named entities recognized.
Entity Linking (EL)
We consider an entity popular if it can be linked to a Wikipedia article. Hence, we run EL for all entities identified in the previous stage to determine if an entity has an associated Wikipedia article that can be linked to. To do this, we make use of a state-of-the-art multilingual EL model, mGENERE (De Cao et al. 2022) for linking entities (either coarse-grained or fine-grained). Given an entity-tagged sentence (the start and end of an entity are indicated explicitly with the special tags [START] and [END]), mGENRE outputs a ranked list of $n$ Wikipedia page titles along with their associated log probabilities, from which we make use of the top item with the highest probability to determine if indeed the entity has a relevant Wikipedia article that can be linked to. If the log probability is above a threshold, we consider that entity as popular.
During parameter tuning, we observed that the variation in threshold did not significantly affect the performance of the claim detection models that utilize EL. This could be due to the majority of named entities being recognized as Wikipedia entities, rendering the information redundant for the model. A possible cause of this nature is NER models using Wikipedia as one of the sources to generate the NER training data. Consequently, we set the EL threshold to -0.15 (probability score of 0.861), as this value resulted in a significant drop in the number of entities linked (refer to Figure 5 in Appendix A), ensuring that only highly probable entity links are used to determine the popularity. Refer to Appendix A for the statistics of named entities linked.
Entity Type Embedding
Similar to the word embedding, an entity type can be represented in a dense vector representation, enabling the model to capture entity type information required for the task as a dense vector. This is achieved by adding an embedding layer to the claim detection model which converts each entity type to an entity type vector. Based on the entity type leveraged, we experiment with the following four variants of our EX-Claim model:
  • EX-Claim-C: Uses four coarse-grained named entities recognized and other indicating words that are not named entities.
  • EX-Claim-CP: Uses both coarse-grained NER and the popularity of an entity. Each named entity can be either popular or unpopular, doubling the size of the entity types. This results in 9 entity types (2*coarse-grained NER types + 1).
  • EX-Claim-F: Uses 15 fine-grained named entities recognized and other indicating words that are not named entities.
  • EX-Claim-FP: Uses both fine-grained NER and the popularity of an entity, resulting in 31 entity types (2*fine-grained NER types + 1).
Category Entity Type Criteria
No of.
Entity Types
X-Claim None 0
EX-Claim-C Coarse-grained NER 5
EX-Claim-CP Coarse-grained NER & Entity Linking 9
EX-Claim-F Fine-grained NER 15
EX-Claim-FP Fine-grained NER & Entity Linking 31
Table 1: Entity Type Categories Experimented
Table 1 summarizes the entity type categories experimented in this study. As an ablated variant, we experiment with X-Claim category which does not use any entity-type information to show the effectiveness of utilizing entity-type information for verifiable claim detection. Except for the X-Claim category, each word in the input sequence is identified with an entity type index $et_{i}$ according to one of the entity type categories. This information is fed into the entity type embedding layer to obtain a dense vector representation $ee_{i}$ for each word in the input sequence.
Entity-aware Claim Detection
Notation Description
$w_{i}$ $i^{th}$ word from the input sequence
$n$ Number of words in the input sequence
$we_{i}\in\mathbb{R}^{d_{w}}$ Word embedding of $i^{th}$ word
$d_{w}$ Dimension of word embedding
$et_{i}\in\mathbb{R}^{k}$ One-hot vector indicating entity type of $i^{th}$ word
$k$ Number of entity types
$ee_{j}\in\mathbb{R}^{d_{e}}$ Entity type embedding of $j^{th}$ entity type
$d_{e}$ Dimension of entity type embedding
$d_{p}$ Projection layer dimension
$P_{w}\in\mathbb{R}^{d_{w}\times d_{p}}$ Word embedding projection matrix
$P_{e}\in\mathbb{R}^{d_{e}\times d_{p}}$ Entity type embedding projection matrix
$h_{i}\in\mathbb{R}^{d_{p}}$ Projected input representation
$H$ Projected input sequence $(h_{1},h_{2},...h_{n})$
$se\in\mathbb{R}^{d_{p}}$ Sentence embedding
$C$ Number of class labels
$y\in\mathbb{R}^{C}$ Probability distribution predicted by the model for the input sequence
$\hat{y}\in\mathbb{R}^{C}$ Ground truth probability distribution of the input sequence
$EE\in\mathbb{R}^{k\times d_{e}}$ Entity embedding layer
$Attn$ Attention layer
Table 2: List of Key Notations
Given an input word sequence $(w_{1},w_{2},\ldots,w_{n})$, we extract the cross-lingual word embeddings $(we_{1},we_{2},\ldots,we_{n})$ , and the entity type index $(et_{1},et_{2},\ldots,et_{n})$. Here, $et_{i}\in\mathbb{R}^{k}$ represents the entity type of a word as a one-hot vector, where $k$ indicates the number of entity types. These entity-type indices are used to extract entity-type embeddings as follows,
$EE(et_{1},et_{2},...et_{n})=(ee_{1},ee_{2},...ee_{n})$ (1)
Once the entity-type embeddings are obtained, we project both word embeddings and entity-type embeddings into a common vector space and add the projected vectors at the word level as follows,
$h_{i}=P_{w}\cdot we_{i}+P_{e}\cdot ee_{i}\in\mathbb{R}^{d_{p}}$ (2)
This projection layer is responsible for converting the input features into task-specific, low-dimensional vectors and incorporating entity-type information into the word embedding.
To attend to the important words indicating the verifiability of claims, we apply scaled dot-product attention (Vaswani et al. 2017) to the projected input sequence $H=(h_{1},h_{2},...h_{n})$ as follows,
$Attn(H)=softmax(\frac{HH^{T}}{\sqrt{d_{p}}})\cdot H$ (3)
The attention layer $Attn$ generates attended context vectors for each element in the input sequence. The mean value of this context vector is obtained as the sentence representation $se$ as follows,
$se=\frac{1}{n}\sum Attn(H)\in\mathbb{R}^{2d_{p}}$ (4)
Finally, the sentence embedding $se$ is fed into a fully connected layer followed by a softmax function to generate a probability distribution $y$ over the output classes as follows,
$out=W_{o}se+b_{o}\in\mathbb{R}^{C}$ (5)
$y_{j}=\frac{exp(out_{j})}{\sum_{j}^{C}exp(out_{j})}\in\mathbb{R}^{C}$ (6)
Given the true label distribution $\hat{y}\in\mathbb{R}^{C}$, the entity type embedding layer, bidirectional GRU, and attention layer are trained by optimizing the cross-entropy loss as follows,
$L_{CE}=-\sum_{j=1}^{C}\hat{y}_{j}log(y_{j})$ (7)
Experiment Setting
Dataset
Detecting claims across languages, and evaluating the model’s capability in identifying the same claim across multiple languages requires a parallel and multilingual ground truth dataset. We use CheckThat! 2022 (Nakov et al. 2022) dataset for training and testing the claim detection models. This dataset contains tweets related to COVID-19, written in Arabic, Bulgarian, Dutch, English and Turkish. Further, it contains four partitions: Train, Dev, Dev-test, and Test. We used the Train partition of all languages except Arabic for training and used the Dev partition as validation data. The Arabic data was excluded during the training as the initial experiments yielded poor performance in all languages including Arabic. Further, the Dev-Test partition was used to find the optimal parameters.
To further evaluate the capability of the models in identifying claims written in a broad range of unseen languages, we generated synthetic data from CheckThat! Test partition. The synthetic data was generated for 18 unseen global languages. For each unseen language, we randomly sampled 250 data instances from the Test partition of CheckThat! data. Sampled data instances were translated to the target language using Microsoft Azure AI translator (Junczys-Dowmunt 2019). Every time, a new set of 250 data instances was sampled to ensure that the same set of tweets did not appear across the languages.
To evaluate the generalization capability of the proposed model in identifying claims discussing unseen topics as well as unseen language, we report the performance in Kazemi 2021 (Kazemi et al. 2021). This dataset contains WhatsApp tipline and social group messages written in in high-resource languages such as English and Hindi, and low-resource languages such as Bengali, Malayalam, and Tamil. Refer to Appendix A for the statistics of the datasets.
The claim detection models were evaluated using these three datasets gathered from different social media platforms, enabling assessment of generalization of the proposed solution across the following scenarios:
  • Known context and seen languages - Models were tested on the CheckThat! 2022 test partition. Except for Arabic, the rest of the languages were seen by the model during the training.
  • Known context and unseen languages - Models were tested on synthetic data composed of 18 languages that were not present in the training data.
  • Unknown context and seen/unseen languages - Models were tested on the Kazemi 2021 dataset. Except for English, the other four languages are unseen during the training.
Hyperparameters
Hyperparameter Value
Projection Size 256
Learning Rate 3e-5
Entity Embedding Size 128 (C,F), 256 (CP,FP)
Entity Linking Threshold -0.15
Table 3: Hyperparameters Used
Table 3 lists the hyperparameters used. Following (Devlin et al. 2019), we set the maximum length of the input sequence to 128, and use Adam as the optimizer. We train the models for 30 epochs and choose the model with the lowest validation loss as the best model. Initial experiments revealed, that the proposed solution performed better when XLM-R was used as the cross-lingual embedding layer instead of mBERT. Therefore, the rest of the experiments use XLM-R as the cross-lingual word embedding layer in EX-Claim.
Model Accuracy Precision Recall F1-Score
CheckThat! 2022
mBERT 0.722 $\pm$ 0.006 0.723 $\pm$ 0.006 0.722 $\pm$ 0.006 0.715 $\pm$ 0.007
XLM-R 0.669 $\pm$ 0.007 0.695 $\pm$ 0.006 0.669 $\pm$ 0.007 0.658 $\pm$ 0.009
mT5 0.681 $\pm$ 0.004 0.685 $\pm$ 0.004 0.681 $\pm$ 0.004 0.679 $\pm$ 0.004
Llama2 0.776 $\pm$ 0.001 0.782 $\pm$ 0.001 0.776 $\pm$ 0.001 0.771 $\pm$ 0.001
Mistral 0.769 $\pm$ 0.003 0.784 $\pm$ 0.003 0.769 $\pm$ 0.003 0.761 $\pm$ 0.003
Phi3 0.744 $\pm$ 0.001 0.75 $\pm$ 0.001 0.744 $\pm$ 0.001 0.738 $\pm$ 0.001
Entity Types Categories X-Claim 0.747 $\pm$ 0.005 0.751 $\pm$ 0.005 0.747 $\pm$ 0.005 0.745 $\pm$ 0.005
EX-Claim-C 0.752 $\pm$ 0.005 0.755 $\pm$ 0.005 0.752 $\pm$ 0.005 0.75 $\pm$ 0.005
EX-Claim-CP 0.753 $\pm$ 0.007 0.756 $\pm$ 0.007 0.753 $\pm$ 0.007 0.751 $\pm$ 0.007
EX-Claim-F 0.754 $\pm$ 0.005 0.758 $\pm$ 0.006 0.754 $\pm$ 0.005 0.752 $\pm$ 0.006
EX-Claim-FP 0.755 $\pm$ 0.007 0.759 $\pm$ 0.007 0.755 $\pm$ 0.007 0.754 $\pm$ 0.007
Synthetic Data
mBERT 0.687 $\pm$ 0.009 0.704 $\pm$ 0.008 0.687 $\pm$ 0.009 0.679 $\pm$ 0.01
XLM-R 0.662 $\pm$ 0.008 0.68 $\pm$ 0.007 0.662 $\pm$ 0.008 0.656 $\pm$ 0.009
mT5 0.659 $\pm$ 0.007 0.662 $\pm$ 0.007 0.659 $\pm$ 0.007 0.658 $\pm$ 0.007
Llama2 0.724 $\pm$ 0.002 0.749 $\pm$ 0.002 0.724 $\pm$ 0.001 0.707 $\pm$ 0.002
Mistral 0.718 $\pm$ 0.002 0.759 $\pm$ 0.003 0.718 $\pm$ 0.002 0.701 $\pm$ 0.003
Phi3 0.691 $\pm$ 0.002 0.735 $\pm$ 0.005 0.691 $\pm$ 0.002 0.664 $\pm$ 0.002
Entity Types Categories X-Claim 0.719 $\pm$ 0.007 0.724 $\pm$ 0.006 0.719 $\pm$ 0.007 0.718 $\pm$ 0.007
EX-Claim-C 0.729 $\pm$ 0.007 0.734 $\pm$ 0.007 0.729 $\pm$ 0.007 0.728 $\pm$ 0.007
EX-Claim-CP 0.727 $\pm$ 0.01 0.733 $\pm$ 0.008 0.727 $\pm$ 0.01 0.727 $\pm$ 0.01
EX-Claim-F 0.73 $\pm$ 0.008 0.735 $\pm$ 0.007 0.73 $\pm$ 0.008 0.729 $\pm$ 0.008
EX-Claim-FP 0.731 $\pm$ 0.009 0.737 $\pm$ 0.008 0.731 $\pm$ 0.009 0.73 $\pm$ 0.009
Kazemi 2021 Data
mBERT 0.716 $\pm$ 0.007 0.725 $\pm$ 0.006 0.716 $\pm$ 0.007 0.702 $\pm$ 0.009
XLM-R 0.687 $\pm$ 0.011 0.752 $\pm$ 0.004 0.687 $\pm$ 0.011 0.686 $\pm$ 0.012
mT5 0.681 $\pm$ 0.004 0.725 $\pm$ 0.004 0.681 $\pm$ 0.004 0.678 $\pm$ 0.005
Llama2 0.633 $\pm$ 0.003 0.5 $\pm$ 0.001 0.633 $\pm$ 0.001 0.613 $\pm$ 0.004
Mistral 0.603 $\pm$ 0.004 0.747 $\pm$ 0.005 0.603 $\pm$ 0.004 0.55 $\pm$ 0.005
Phi3 0.546 $\pm$ 0.001 0.744 $\pm$ 0.001 0.546 $\pm$ 0.001 0.455 $\pm$ 0.001
Entity Types Categories X-Claim 0.788 $\pm$ 0.007 0.817 $\pm$ 0.005 0.788 $\pm$ 0.007 0.789 $\pm$ 0.007
EX-Claim-C 0.786 $\pm$ 0.009 0.806 $\pm$ 0.007 0.786 $\pm$ 0.009 0.789 $\pm$ 0.008
EX-Claim-CP 0.785 $\pm$ 0.007 0.803 $\pm$ 0.007 0.785 $\pm$ 0.007 0.788 $\pm$ 0.007
EX-Claim-F 0.793 $\pm$ 0.008 0.813 $\pm$ 0.006 0.793 $\pm$ 0.008 0.796 $\pm$ 0.008
EX-Claim-FP 0.794 $\pm$ 0.008 0.813 $\pm$ 0.006 0.794 $\pm$ 0.008 0.797 $\pm$ 0.007
All
mBERT 0.698 $\pm$ 0.008 0.711 $\pm$ 0.007 0.698 $\pm$ 0.008 0.689 $\pm$ 0.009
XLMR 0.668 $\pm$ 0.008 0.695 $\pm$ 0.006 0.668 $\pm$ 0.008 0.661 $\pm$ 0.01
mT5 0.667 $\pm$ 0.006 0.677 $\pm$ 0.006 0.667 $\pm$ 0.006 0.665 $\pm$ 0.006
Llama2 0.717 $\pm$ 0.002 0.755 $\pm$ 0.001 0.717 $\pm$ 0.002 0.702 $\pm$ 0.002
Mistral 0.707 $\pm$ 0.003 0.761 $\pm$ 0.004 0.707 $\pm$ 0.003 0.685 $\pm$ 0.003
Phi3 0.674 $\pm$ 0.002 0.74 $\pm$ 0.004 0.674 $\pm$ 0.002 0.64 $\pm$ 0.002
Entity Types Categories X-Claim 0.736 $\pm$ 0.006 0.746 $\pm$ 0.006 0.736 $\pm$ 0.006 0.735 $\pm$ 0.007
EX-Claim-C 0.743 $\pm$ 0.007 0.751 $\pm$ 0.006 0.743 $\pm$ 0.007 0.743 $\pm$ 0.007
EX-Claim-CP 0.742 $\pm$ 0.009 0.75 $\pm$ 0.008 0.742 $\pm$ 0.009 0.742 $\pm$ 0.009
EX-Claim-F 0.746 $\pm$ 0.007 0.753 $\pm$ 0.007 0.746 $\pm$ 0.007 0.745 $\pm$ 0.008
EX-Claim-FP 0.747 $\pm$ 0.008 0.754 $\pm$ 0.007 0.747 $\pm$ 0.008 0.746 $\pm$ 0.009
Table 4: Dataset-level Performance of the Claim Detection Models
Evaluation metrics
We report the accuracy, precision, recall, and F1-score of the claim detection models. We fine-tuned ten models and report the average and standard deviations for robust comparison. Since each language is composed of different ratios of verifiable and unverifiable claims, we computed the weighted precision, recall, and F1-score at the language level, and the overall performance is obtained by averaging the language-level performance of the models.
Baseline models
In the absence of cross-lingual claim detection models in the literature, we leverage strong cross-lingual language models as our baselines. We compare mBERT (Alam et al. 2021b; Uyangodage, Ranasinghe, and Hettiarachchi 2021; Panda and Levitan 2021; Zengin, Kartal, and Kutlu 2021; Hasanain and Elsayed 2022), XLM-RoBERTa (XLM-R) (Alam et al. 2021b, a), and mT5 (Du, Gollapalli, and Ng 2022) fine-tuned on the CheckThat! Train partition as the baselines. These models are widely used as a solution to handle multilingual claims and have shown to be consistently competitive across claim detection studies. In addition to these transformer-based language models, we compare some of the open-sourced large language models experimented in Li, Panchendrarajan, and Zubiaga (2024) as the baselines. This includes Llama2-7B, Mistral-7B, and Phi3 fine-tuned on the same training data for the text generation task. Refer to Appendix C for more details on the prompt used and the LLM fine-tuning process.
Results
Baseline Entity Type Categories
Language
mBERT
XLM-R
mT5
Llama2
Mistral
Phi3
X-Claim
EX-Claim-C
EX-Claim-CP
EX-Claim-F
EX-Claim-FP
Arabic 0.692 0.654 0.677 0.752 0.736 0.718 0.731 0.733 0.736 0.733 0.735
Bengali 0.725 0.655 0.556 0.697 0.642 0.552 0.713 0.759 0.755 0.748 0.764
Bulgarian 0.783 0.633 0.729 0.835 0.815 0.812 0.782 0.787 0.789 0.79 0.796
Czech 0.707 0.647 0.72 0.808 0.786 0.755 0.741 0.778 0.775 0.774 0.769
Dutch 0.68 0.698 0.634 0.72 0.657 0.663 0.716 0.721 0.72 0.717 0.718
English (Kaz) 0.738 0.769 0.748 0.764 0.75 0.719 0.755 0.747 0.743 0.757 0.755
English (CT) 0.724 0.658 0.688 0.769 0.777 0.755 0.75 0.758 0.762 0.762 0.761
French 0.638 0.698 0.658 0.736 0.739 0.738 0.703 0.724 0.724 0.722 0.725
German 0.751 0.714 0.722 0.808 0.822 0.816 0.789 0.77 0.767 0.778 0.775
Greek 0.677 0.64 0.71 0.697 0.687 0.679 0.735 0.748 0.746 0.747 0.759
Gujarati 0.661 0.669 0.649 0.457 0.582 0.455 0.771 0.775 0.763 0.77 0.751
Hindi 0.631 0.695 0.613 0.792 0.809 0.6 0.766 0.761 0.751 0.767 0.769
Hungarian 0.667 0.651 0.665 0.757 0.727 0.637 0.68 0.682 0.674 0.683 0.682
Italian 0.721 0.647 0.704 0.764 0.745 0.736 0.709 0.724 0.72 0.731 0.723
Japanese 0.716 0.658 0.693 0.759 0.771 0.74 0.715 0.728 0.728 0.732 0.746
Kannada 0.535 0.567 0.579 0.371 0.459 0.342 0.628 0.666 0.674 0.674 0.676
Korean 0.691 0.635 0.623 0.69 0.72 0.653 0.712 0.724 0.724 0.718 0.709
Marathi 0.588 0.64 0.603 0.643 0.532 0.46 0.668 0.664 0.657 0.666 0.664
Malayalam 0.79 0.564 0.685 0.374 0.151 0.141 0.845 0.826 0.832 0.84 0.835
Persian 0.673 0.669 0.622 0.713 0.667 0.602 0.704 0.718 0.717 0.716 0.718
Russian 0.709 0.667 0.641 0.766 0.764 0.734 0.715 0.721 0.718 0.724 0.724
Spanish 0.738 0.693 0.683 0.779 0.735 0.77 0.787 0.783 0.78 0.783 0.784
Swedish 0.707 0.688 0.654 0.772 0.779 0.76 0.777 0.775 0.776 0.778 0.787
Tamil 0.623 0.744 0.789 0.439 0.398 0.264 0.864 0.853 0.86 0.866 0.861
Thai 0.628 0.648 0.644 0.716 0.698 0.673 0.726 0.711 0.709 0.716 0.715
Turkish 0.699 0.648 0.664 0.775 0.82 0.745 0.748 0.751 0.749 0.759 0.758
Ukrainian 0.688 0.597 0.604 0.764 0.724 0.719 0.644 0.692 0.691 0.688 0.707
Vietnamese 0.727 0.672 0.676 0.735 0.694 0.689 0.715 0.714 0.736 0.717 0.729
Table 5: F1-Score of the Claim Detection Models at Language-level
Dataset-level Performance
Table 4 presents the performance of the claim detection models. The results demonstrate that EX-Claim significantly outperforms the baseline models across all four metrics in overall evaluation. This underscores the importance of learning task-specific representations and attending to critical features to identify the verifiability of a claim. Furthermore, the consistent performance across diverse datasets from various platforms covering multiple topics suggests that our hypothesis is generalizable, regardless of the nature of the claim. It can be observed that the large language models, Llama2 and Mistral achieve the highest performance in CheckThat! dataset. However, their performance drops significantly on the other two datasets, indicating challenges in handling a diverse range of languages. Among the transformer-based baselines, mBERT model outperforms the other two across all datasets. However, utilizing mBERT as the cross-lingual embedding layer in EX-Claim did not yield the optimal performance. This shows that mBERT is powerful in representing cross-lingual context vectors, whereas, XLM-R is good at generating cross-lingual word representations.
Moreover, incorporating entity-type information always enhances the accuracy, precision, recall, and F1-score in CheckThat! and synthetic datasets. This increment is significant in the synthetic dataset composed of 18 unseen languages with at least a 1.2% increment across all the metrics. This shows that, while knowing the context, the EX-Claim utilizes entity information to improve the performance of both seen and unseen languages. Among these, the EX-Claim-FP model which uses fine-grained NER information and EL stands out as the best model across all three datasets. It is interesting to note that the NER models (EX-Claim-C or EX-Claim-F) and their corresponding models utilizing EL information (EX-Claim-FP, EX-Claim-CP) reach very similar performances. This similarity suggests that most of the entities identified as NER are already recognized as popular, resulting in only minor performance improvements.
Language-level Performance
Table 5 lists the language-level F1-score of the claim detection models across 27 languages present in the test datasets. It is noteworthy that large language models (LLMs) such as Llama2 and Mistral achieve the highest performance across high-resource languages. However, their performance declines sharply in various low-resource languages. On the contrary, EX-Claim not only outperforms LLMs in several low-resource languages but also delivers competitive performance in high-resource languages. This highlights that transformer-based models like XLM-R remain highly competitive for text classification tasks, despite their significantly smaller size compared to larger models (e.g., 250M versus 7B parameters). Furthermore, consistent with the overall trends discussed earlier, at least one (although generally all) of the variants of the EX-Claim significantly improves the transformer-based baselines across all the languages except in the English language in the Kazemi dataset. Further, the models that leverage entity information show notable improvements in several languages, with particularly significant gains observed in Bengali, Czech, French, Greek, Italian, Japanese, Kannada, and Ukrainian. However, it is noteworthy that the inclusion of entity information results in a performance decline in certain languages, such as German, Malayalam, and Thai.
Analysis
Cross-lingual Knowledge Transfer
The synthetic data comprises claims from the CheckThat! Test partition, written in 18 different languages. We use this dataset to assess the cross-lingual knowledge-transferring capabilities of the claim detection models. We compute the fraction of data instances in the Synthetic data which gets the same class prediction as its original data instance from the CheckThat! Test partition by a model. We categorized these results into correct and incorrect predictions, allowing us to evaluate the model’s ability to transfer knowledge across correct and wrong predictions.
Model
Correct
Prediction
Wrong
Prediction
mBERT 82.75% 64.83%
XLMR 81.72% 66.27%
mT5 79.12% 60%
Llama2 85.4% 70.1%
Mistral 86.87% 71.3%
Phi3 82.46% 65.5%
Entity Types Categories X-Claim 87.21% 71.73%
EX-Claim-C 88.15% 68.76%
EX-Claim-CP 87.43% 68.63%
EX-Claim-F 87.82% 68.48%
EX-Claim-FP 87.09% 66.51%
Table 6: Cross-lingual Knowledge Transfer Rate
Table 6 lists the knowledge transfer rate of claim detection models. Here, the transformer-based baseline models exhibit the lowest transfer rate in correct and wrong predictions. In contrast, EX-Claim significantly enhances the transfer rate among the correct and wrong predictions and outperforms the LLMs compared when transferring knowledge across languages in correct prediction. Further, the incorporation of entity information tends to either maintain or improve the transfer rate of correct predictions, while consistently reducing the transfer rate of incorrect predictions. This behavior indicates that the projection layer, which maps input features into a task-specific lower-dimensional vector, effectively transfers knowledge between languages. Additionally, the entity information enhances language-level performance, thereby reducing the transfer rate of incorrect predictions.
Attending to Important Words
The attention layer enables the claim detection models to attend to important words in the input sequence before generating the sentence representation. We quantify this behavior of the models as the entropy of attention weights computed by the attention layer. We compute the average entropy of the test data by computing the mean value of entropy of all the input sequences.
Model Entropy
X-Claim 4.09
EX-Claim-C 3.579
EX-Claim-CP 3.479
EX-Claim-F 3.519
EX-Claim-FP 3.322
Table 7: Entropy of Attention Layer
Table 7 lists the attention layer’s entropy of the claim detection models. We can see that the models utilizing entity information always yield lower entropy. This shows that these models are concentrated on a few keywords in the input sequence.222 Refer to Appendix B for attention weights generated for sample sentences. Moreover, the entity types identified via the popularity further reduce the entropy of the models.
Entity-type Embedding
(a) EX-Claim-CP
(b) EX-Claim-F
Figure 2: Two-dimensional embedding visualization of EX-Claim-CP and EX-Claim-F Models
Unlike the cross-lingual word embedding which is directly obtained from the pre-trained models, EX-Claim learns a dense vector representation for each entity type. Figures 2(a) and 2(b) show the 2D visualization of the entity type embeddings learned by the EX-Claim-CP and EX-Claim-F models. Popular entities are indicated with the post-fix -P. The 2D vector representation of the embeddings is obtained using the t-SNE [^3] https://scikit-learn.org/stable/modules/generated/sklearn
.manifold.TSNE.html
dimension reduction technique. Further, the embedding representations were clustered using the KMeans clustering technique (Sinaga and Yang 2020), and the number of optimal clusters was chosen using Silhouette score (Rousseeuw 1987).
The EX-Claim-CP model learns three groups of entity types: the location entities form one cluster, both popular and unpopular miscellaneous entities form another cluster, and the remaining entities, irrespective of their popularity, constitute the third cluster. Similarly, EX-Claim-F model learns various groups of entity types. Notably, this model clusters Animal, Disease, Person, Biological Entity, Disease, Plan and Mythological Entity into a single group, likely influenced by the COVID-19 discussions present in the training data. This observation highlights the model’s tendency to learn context-specific entity-type embeddings during training.
Ablation Study
We conduct an ablation study to identify the impact of the key components of EX-Claim with the following variations.
  • No projection layer (NPL) - The projection layer is omitted, and attention is applied directly to the embedding representations.
  • No attention layer (NAL) - The attention layer is removed, hence the projected vector representations are aggregated into a sentence embedding by computing their mean.
  • No embedding layer (NEL) - The embedding layer is removed, and the entity information is represented as one-hot vectors.
Model F1-Score
X-Claim 0.735
X-Claim - NPL 0.688
X-Claim - NAL 0.714
Table 8: Ablation Results of X-Claim Model
The NLP and NAL variations are applied to X-Claim, and Table 8 presents the overall F1-score of these models. The results demonstrate that the projection layer is crucial for learning task-specific input representations, as its removal leads to a 4.7% drop in overall performance. Likewise, the attention layer enhances the model’s performance, contributing to a 2.1% improvement.
Model F1-Score
EX-Claim-C 0.743
EX-Claim-C - NEL 0.732
EX-Claim-CP 0.742
EX-Claim-CP - NEL 0.731
EX-Claim-F 0.745
EX-Claim-F - NEL 0.733
EX-Claim-FP 0.746
EX-Claim-FP - NEL 0.733
Table 9: Ablation Results of EX-Claim Models
Table 9 lists the performance of EX-Claim models with NEL variation. It can be seen, that the performance declines by approximately 1.1% - 1.3% when entity information is represented as one-hot vectors, suggesting that the models struggle to effectively utilize entity information when it is provided as a one-hot vector representation.
Error Analysis
False Positive
S4 Months after receiving the second dose of the Pfizer COVID vaccine, I continue to suffer side effects including perfect health and having to deal with morons.
S5 How true is Congress party’s claim regarding shortfall in Government earnings
False Negative
S6 These are not photos of a man with an artificial heart
S7 Retirement Home NOT Raided By The FBI For Running Elderly Fight Club, NO 7 Arrests
S8 Gruesome Video Of A Man Hacked To Death In Bihar Falsely Shared As TMC ’Goons’ Attacking A BJP Worker
S9 Video Showing Suicide Bid Falsely Linked To CAA
Table 10: Misclassified Sample Sentences
We analyze the false positive and false negative predictions made by all four variants of EX-Claim. Table 10 lists some of the misclassified examples. We noticed that the appearance of disease names ($S4$), lengthy claims, nonrecognition of personal experience from pronouns’ presence ($S4$), and ambiguous fact expressions could cause false positive scenarios ($S5$).
We observed that misclassification of popular events expressing personal experience tends to be a common cause of false negatives. Refer to Sentences S6-S9 in Table 10 for sample false negative instances where the reach of the event makes it verifiable, yet the personal experience framing leads to misclassification. Further, some of these claims refer to other modalities such as images and videos in the expression of facts, making it further difficult when relying solely on textual information  [^4] Refer to Appendix B showing the correlation between the confusion matrix and some of these features..
Conclusion
We present EX-Claim, the first cross-lingual claim detection model that encapsulates entity information, including their popularity, for improved performance in detecting claims spreading on social media, across multiple languages. Extensive experiments on three datasets comprising 27 languages reveal that EX-Claim significantly outperforms competitive baseline methods. Especially EX-Claim shows competitive performance across high-resource languages and performs better in low-resource languages than the large language models compared. Further, the language-level analysis shows that certain languages particularly benefit from entity information, resulting in improved performance. Furthermore, the models achieve the highest cross-lingual knowledge transfer rate by accurately transferring knowledge from correct predictions while minimizing errors during transfer, a critical capability for addressing misinformation on diverse social media platforms. Our in-depth analysis reveals that EX-Claim achieves the lowest average entropy in attention weights, indicating a focus on key terms to assess verifiability. Future work will explore techniques to generalize entity-type embeddings beyond the context encountered in the training data and address the challenges contributing to false positive and false negative outcomes identified in our error analysis.
Limitations
The limitations of this work are as follows:
  • Training data limited to a single topic - Our training data consists solely of COVID-19 discussions. Although our model generalizes well across different datasets, including those focused on political topics, this limitation may still restrict the model’s broader generalization capabilities.
  • Dependence on NER and entity linking tools: The proposed solution relies on existing cross-lingual NER and entity linking tools to extract entity-type information. The performance of these tools directly impacts the overall effectiveness of our model, particularly in languages where NER or entity-linking tools may perform poorly or exceptionally well.
  • Training data limited to a single platform: Our training data is exclusively composed of tweets, which may cause the model to learn characteristics specific to this platform. While we demonstrate that our model can handle data from other platforms, such as WhatsApp tiplines in the Kazemi dataset, this limitation could still hinder the model’s ability to generalize across different types of social media content.
  • Exclusion of popular events not identified by a name: Since the system identifies and tracks the popularity of named entities, popular events not linked to any names are excluded. As previously discussed, this is a significant cause of false negative predictions.
Acknowledgments
This project is funded by the European Union and UK Research and Innovation under Grant No. 101073351 as part of Marie Skłodowska-Curie Actions (MSCA Hybrid Intelligence to monitor, promote, and analyze transformations in good democracy practices).
Appendix A Statistics
Statistics of Named Entity Types Identified
Table 11 lists the types of named entities identified using the coarse-grained and fine-grained NER models. As discussed earlier, WikiNEuRal identifies 4 types of coarse-grained named entities and MultiNERD identifies 15 types of fine-grained Named Entities.
NER Model Entities
WikiNEuRal
Person (PER)
Organization (ORG)
Location (LOC)
Miscellaneous (MISC)
MultiNERD
Person (PER)
Organization (ORG)
Location (LOC)
Animal (ANIM)
Biological entity (BIO)
Celestial Body (CEL)
Disease (DIS)
Event (EVE)
Food (FOOD)
Instrument (INST)
Media (MEDIA)
Plant (PLANT)
Mythological entity (MYTH)
Time (TIME)
Vehicle (VEHI)
Table 11: Named Entity types Identified by the NER Models
Figure 3: Average Number of Fine-grained Entities Identified in the Train Partition of CheckThat! 2022 Data
Figure 4: Average Number of Coarse-grained Entities Identified in the Train Partition of CheckThat! Data
Statistics of Named Entities Recognized
Figures 4 and 3 show the average number of coarse-grained and fine-grained named entities recognized respectively in the Train partition of CheckThat! 2022 data. It can be observed that the average number of coarse-grained location entities present in a sentence is relatively higher for all languages compared to the other three types. This could be due to the COVID-19 discussion in the training data referring to the status of the disease from different countries and locations. Miscellaneous entities stand out as the 2nd most frequent coarse-grained entity type for all languages except for Arabic. When comparing coarse-grained and fine-grained NERs, the distribution across languages is similar for the organization entity type. However, both tools recognize different distributions for person and location entity types. Since the dataset contains tweets about COVID-19, the occurrences of disease, media, biological entities, and food are relatively higher compared to other rare entity types.
Statistics of Entities Linked
(a) Coarse-grained EL
(b) Fine-grained EL
Figure 5: Number of Entities Linked with the Variation in EL Threshold in Train Partition of CheckThat! Data
(a) Popular vs Unpopular Coarse-grained Entities
(b) Popular vs Unpopular Fine-grained Entities
Figure 6: Fraction of Popular and Unpopular Entities in Train Partition of CheckThat! Data
Figure 5 shows the number of named entities linked to a Wikipedia page with the increase in threshold. It can be observed that the reduction in the number of entities linked to a Wikipedia page with the increase in log probability threshold is not very significant until the threshold is increased to -0.15 (probability score of 0.861). This indicates that most of the coarse-grained or fine-grained NER identified are linked to a Wikipedia page with a very high probability.
Figure 6 shows the fraction of popular and unpopular entities determined with the log probability threshold of -0.15 (probability of 0.861). Both diagram shows a similar pattern, where Arabic language data contains more fraction of entities identified as popular followed by English, and Dutch and Turkish have the lowest fraction among all. Further, the chances of fine-grained entities being popular entities are slightly higher (ranging from 0.39 to 0.63) compared to coarse-grained entities (ranging from 0.33 to 0.61). Moreover, this indicates, that nearly more than 40% of the entities recognized are linked as popular ones in both categories with a very high probability.
Statistics of the Datasets Used
We use two datasets that are publicly available for the experiments, and the datasets may contain personally identifiable information or offensive content.
CheckThat! Dataset
Arabic Bulgarian Dutch English Turkish Total
Train Verifiable 2,513 1,871 929 2,122 1,589 14,032
Unverifiable 1,118 839 1,021 1,202 828
Total 3,631 2,710 1,950 3,324 2,417
Dev Verifiable 235 177 72 195 150 1,300
Unverifiable 104 74 109 112 72
Total 339 251 181 307 222
Dev-Test Verifiable 691 519 252 574 438 3,837
Unverifiable 305 217 282 337 222
Total 996 736 534 911 660
Test Verifiable 682 199 608 149 303 3,698
Unverifiable 566 130 750 102 209
Total 1,248 329 1,358 251 512
Table 12: Statistics of the CheckThat! 2022 Verifiable Claim Detection Dataset
Table 12 presents the statistics of the CheckThat! 2022 dataset. Compared to other languages, the number of data instances is low in Dutch across all partitions except Test. Further, the ratio between verifiable and unverifiable claims is equal for Dutch languages, whereas other languages contain more verifiable statements than unverifiable (roughly 2:1 ratio).
Kazemi Dataset
Bengali English Hindi Malayalam Tamil Total
Verifiable 329 535 407 707 197 4,017
Unverifiable 593 388 465 263 133
Total 922 923 872 970 330
Table 13: Statistics of the Kazemi 2021 Verifiable Claim Detection Dataset
Table 13 presents the statistics of the Kazemi dataset. The dataset contains WhatsApp tiplines and social media posts about topics such as politics and COVID-19.
Appendix B Analysis
Attention Weights for Sample Sentences
Figure 8 shows the heatmaps of the attention weights computed by different claim detection models for the following three sentences.
  • S10 - Old video of cop beating woman in Gawalior shared as Delhi police brutality
  • S11 - Donald Trump fund raising email takes CNN anchor’s comments out of context
  • S12 - Trump withdraws US from the Iran nuclear deal
It can be observed, that EX-Claim attends to important words in the sentences, whereas X-Claim generates attention weights distributed across most of the words. Further, an interesting pattern of named entities attending to themselves or other named entities more compared to other words can be observed across all three sentences.
Correlation between Confusion Matrix and Selected Features
Figure 7: Correlation of Sample Features vs Confusion Matrix
Figure 7 presents the correlation scores between the true positive, true negative, false positive, and false negative values and selected features. It can be observed that the length of the claim is positively correlated with true positive and false positive rates, showing the models tend to assume lengthy claims as verifiable.
Similarly, the number of disease names appearing in a claim is positively correlated with a false positive rate. This could be the COVID-19 discussion in the training data making the model to classify the claims with disease names as verifiable. Further, the number of media names is positively correlated with false negative rate as discussed in Section Error Analysis. This suggests that references to media, such as images and videos, can either enhance or hinder verifiability, depending on the context and reach.
(a) S10 : EX-Claim-CP
(b) S10 : X-Claim
(c) S11 : EX-Claim-F
(d) S11 : X-Claim
(e) S12 : EX-Claim-FP
(f) S12 : X-Claim
Figure 8: Heatmap of Attention Weights Generated by Claim Detection Models for sentences S10, S11, and S12
Appendix C Implementation
We fine-tuned the mBERT [^5] https://huggingface.co/google-bert/bert-base-multilingual-cased, XLM-R [^6] https://huggingface.co/FacebookAI/xlm-roberta-base, and mT5 models published in HuggingFace to obtain the transformer-based baseline models. Especially, mT5 is available in various sizes. We used mT5-large [^7] https://huggingface.co/google/mt5-large as this model was the largest model among the variations which we could accommodate with the same resources utilized for training and testing other models. Similarly, the open-source large language models (LLMs) available in HuggingFace were used to obtain the LLM baselines, Llama2888https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, Mistral 7B [^9] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2, and Phi3101010https://huggingface.co/microsoft/Phi-3-mini-4k-instruct. We used the same experimental setting of (Li, Panchendrarajan, and Zubiaga 2024). The prompt 1 used to fine-tune LLMs is a modified version of (Li, Panchendrarajan, and Zubiaga 2024).
### Instruction:
Evaluate whether the input text contains
information or claims that can be verified
through fact-checking. If the statement
presents assertions, facts, or claims,
respond with ’Yes’. If the statement is
purely opinion-based, trivial, or does not
contain any verifiable information or claims,
respond with ’No’.
### Input Sentence: <input sentence>
### Response: <Yes/No>
Prompt 1 Verifiable Claim Detection
All the experiments were conducted using Queen Mary’s Apocrita HPC facility, supported by QMUL Research-IT (King, Butcher, and Zalewski 2017). Specifically, 1 GPU (Volta V100 or Ampere A100) with 8 CPU cores, each composed of 11 GB memory was used to train and test all the models. All the models were fine-tuned by extracting the vector representation of the special token from the final layer and applying a fully connected layer on it for classification. We used only the encoder of the mT5 model for obtaining vector representation of input words. For Named Entity Recognition (NER), we used the WikiNEuRal  [^11] https://huggingface.co/Babelscape/wikineural-multilingual-ner model published in HuggingFace and the resources published for MultiNERD [^12] https://github.com/Babelscape/multinerd/tree/master for fine-tuning the fine-grained NER model. Similarly, the mGENRE [^13] https://huggingface.co/facebook/mgenre-wiki model published in HuggingFace was used for Entity Linking (EL). License of the models used in the experiments can be found at the links provided.