### v3

# Entity-aware Cross-lingual Claim Detection for Automated Fact-checking

Abstract
Identifying claims requiring verification is a critical task in automated fact-checking, especially given the proliferation of misinformation on social media platforms. Despite notable progress, challenges remain—particularly in handling multilingual data prevalent in online discourse. Recent efforts have focused on fine-tuning pre-trained multilingual language models to address this. While these models can handle multiple languages, their ability to effectively transfer cross-lingual knowledge for detecting claims spreading on social media remains under-explored. In this paper, we introduce EX-Claim, an entity-aware cross-lingual claim detection model that generalizes well to handle multilingual claims. The model leverages entity information derived from named entity recognition and entity linking techniques to improve the language-level performance of both seen and unseen languages during training. Extensive experiments conducted on three datasets from different social media platforms demonstrate that our proposed model stands out as an effective solution, demonstrating consistent performance gains across 27 languages and robust knowledge transfer between languages seen and unseen during training.
Entity-aware Cross-lingual Claim Detection for Automated Fact-checking
Rrubaa Panchendrarajan, Arkaitz Zubiaga School of Electronic Engineering and Computer Science Queen Mary University of London {r.panchendrarajan, a.zubiaga}@qmul.ac.uk
1 Introduction
Automated fact-checking is an emerging research task dedicated to combating misinformation, particularly prevalent on social media, and comprises several key stages: claim detection, claim prioritization, evidence retrieval, and claim validation (Zeng et al., 2021). While research in the task is progressing rapidly, identifying and validating claims related to global concerns requires a fact-checking pipeline capable of processing claims written in multiple languages. However, developing multilingual solutions for fact-checking research is not straightforward due to the availability of limited training data, especially for low-resource languages (Panchendrarajan and Zubiaga, 2024). This necessitates advancements in multilingual fact-checking research to leverage limited multilingual training data for developing generalized solutions that can effectively transfer knowledge between languages.
This research focuses on the first component of the fact-checking pipeline: verifiable claim detection. A verifiable claim is defined as a statement expressing facts, excluding personal experience and private knowledge (Panchendrarajan and Zubiaga, 2024). Existing works on verifiable claims detection are predominantly focused on monolingual solutions (Prabhakar et al., 2020; Suri and Dudeja, 2022; Henia et al., 2021; Hussein et al., 2021), with limited attention given to addressing the multilingual nature of the problem. Especially, fine-tuning multilingual language models such as mBERT (Alam et al., 2021b; Uyangodage et al., 2021; Panda and Levitan, 2021) and XLM-R (Alam et al., 2021b; Hüsünbeyi et al., 2022; Savchev, 2022; Alam et al., 2021a) is observed as a common solution. While these models can handle multiple languages, none of them adequately validate the cross-lingual knowledge-transferring capability of these models, particularly in identifying claims written in languages that were not seen during training. To address this challenge, we propose EX-Claim, a cross-lingual claim detection model that can effectively identify verifiable claims regardless of their language. Our model broadens the ability to perform across a diverse set of languages beyond those seen in training by incorporating entity-centric information for knowledge transfer. Our approach leverages the limited multilingual data available in claim detection research to develop a robust cross-lingual model that generalizes beyond the languages seen in the training phase.
We build our model on the assumption that factual claims circulating in social media often revolve around entities. In particular, entity types and their statuses such as their popularity may affect the verifiability of a claim. For instance, consider the following three statements.
  • S1: X visited China to attend EMNLP
  • S2: Keir Starmer visited France to attend Euro Con
  • S3: Keir Starmer visited his home in the UK
Statement $S1$ can be deemed unverifiable from the perspective of a fact-checker, as it describes a personal experience of an unknown individual X. However, a similar context of attending an event expressed in $S2$ is verifiable as the popular entity Keir Starmer (the UK Prime Minister at the time of writing) attending an official event Euro Con becomes verifiable. However, changing the popular entities Keir Starmer or Euro Con with another popular person or event entity will not change the verifiability status of this claim. At the same time, changing the event with a private location such as home reduces the verifiability of the claim $S3$ even if a popular entity is involved. This motivates our intuition that entities, their types, their popularity, and the relationship between them affect the verifiability status of the claims.
Our proposed EX-Claim builds on this intuition by leveraging entity information to effectively identify verifiable claims in a cross-lingual environment. Our model incorporates entity information such as type and popularity derived from named entity recognition (NER) and entity linking (EL). Extensive experiments conducted using three datasets from different social media platforms comprising 27 languages show that EX-Claim stands out as the effective model with consistent performance gain across multiple languages and robust knowledge transfer between languages seen and unseen in training. We make the following key contributions:
  • We propose EX-Claim, a cross-lingual claim detection model for identifying verifiable claims from social media which can effectively transfer knowledge across languages seen and unseen during training.
  • We enhance language-level performance by leveraging entity information in claims through NER and EL.
  • We conduct extensive experiments to evaluate EX-Claim across three datasets representing different social media platforms and encompassing 27 languages, including synthetic data created via machine translation.
The source code and the synthetic data are available online [^1] https://github.com/RubaP/Ex-Claim.
2 Related Work
A common approach to identifying multilingual claims shared on social media involves fine-tuning pre-trained multilingual language models such as mBERT (Alam et al., 2021b; Uyangodage et al., 2021; Panda and Levitan, 2021; Zengin et al., 2021; Hasanain and Elsayed, 2022) and XLM-R (Alam et al., 2021b, a) and developing language-specific models to achieve optimal performance Hüsünbeyi et al. (2022); Savchev (2022); Eyuboglu et al. (2023). Most of these works (Alam et al., 2021b, a; Uyangodage et al., 2021; Baris-Schlicht et al., 2021; Du et al., 2022) fine-tune the language model using combined training data of multiple languages and evaluate the capability of the model in identifying claims written in the same set of languages. Very few studies (Panda and Levitan, 2021; Zengin et al., 2021; Hasanain and Elsayed, 2022) explore the model’s ability to transfer knowledge across language pairs, and these studies are limited to a narrow selection of languages.
Figure 1: EX-Claim : Entity-aware Cross-lingual Claim Detection
With the recent attention to large language models (LLM), Agrestia et al. (2022) utilized GPT-3 to fine-tune the model in each language to develop a language-specific claim detection model. Interestingly, the results revealed that transformer-based architectures like BERT remain competitive with, and in some cases outperform the large language models (LLMs). Following these findings, further research (Sawiński et al., 2023; Lewoniewski et al., 2024; Li et al., 2024) explored leveraging LLMs for claim detection by modeling the task as a text generation problem. While these models demonstrate superior performance in the English language, their ability to handle multilingual claims for the claim detection task has yet to be explored.
Beyond directly using multilingual language models, various methods have been applied to tackle limited training data and noisy social media text. This includes the extraction of platform-specific textual features (Alam et al., 2021b), data augmentation techniques such as machine translation (Zengin et al., 2021; Savchev, 2022) and sampling (Zengin et al., 2021) and multitasking Baris-Schlicht et al. (2021); Du et al. (2022). While these techniques are explored in multilingual settings, none of them have been shown to improve cross-lingual knowledge-transferring capabilities, which our study performs comprehensively for the first time looking at 27 different languages through the exploration of entity-based knowledge transfer.
3 Methodology
Given a statement written in any language, we classify it as verifiable or non-verifiable. Figure 1 illustrates the proposed solution, EX-Claim: entity-aware cross-lingual claim detection framework, whose components are discussed next.
3.1 Cross-lingual Word Embedding
We use XLM-R Kalyan et al. (2021) to extract the cross-lingual vector representation of the words. XLM-R was trained on CommonCrawl data supporting 100 languages, and generates embeddings of size 768 for each tokenized word.
3.2 Named Entity Recognition (NER)
Identifying entities mentioned in the text is essential to extract further information such as its type and popularity. We use state-of-the-art multilingual NER model MultiNERD Tedeschi and Navigli (2022) for identifying named entities. As the model wasn’t released by the authors, we fine-tuned the MultiNERD model with recommended parameters using their source code and training data. MultiNERD recognizes 15 types of fine-grained named entities. Refer to Appendix A.1 for the named entity types and NER statistics.
3.3 Entity Linking (EL)
We consider an entity popular if it can be linked to a Wikipedia article. Hence, we run EL for all entities identified in the previous stage to determine if an entity has an associated Wikipedia article that can be linked to. To do this, we make use of a state-of-the-art multilingual EL model, mGENERE De Cao et al. (2022) for linking entities. Given an entity-tagged sentence, mGENERE outputs a ranked list of $n$ Wikipedia page titles along with their associated log probabilities, from which we make use of the top item with the highest probability to determine if indeed the entity has a relevant Wikipedia article that can be linked to. If the log probability is above a threshold, we consider that entity as popular.
During parameter tuning, we observed that the variation in threshold did not significantly affect the performance of the claim detection models that utilize EL. This is likely because most named entities were already recognized as Wikipedia entities, making the additional information redundant—possibly due to the NER model’s training data being partially sourced from Wikipedia. Consequently, we set the EL threshold to -0.15 (probability 0.861), which notably reduced the number of linked entities (refer to Figure 4 in Appendix A.2), ensuring that only highly probable entity links are used to determine the popularity. Refer to Appendix A.2 for the EL statistics.
3.4 Entity Type Embedding
Similar to the word embedding, an entity type can be represented in a dense vector representation, enabling the model to capture entity type information required for the task as a dense vector. This is achieved by adding an embedding layer to the claim detection model, which converts each entity type to an entity type vector. Based on the entity type leveraged, we experiment with the following two variants of EX-Claim:
  • EXN-Claim: Uses 15 named entity types recognized by the NER model, along with the other category for non-entities—resulting in a total of 16 entity types.
  • EXP-Claim: Uses both NER and the popularity of an entity. Each named entity can be either popular or unpopular, doubling the size of the entity types. This results in 31 entity types (2*15 + 1).
As an ablated variant, we experiment with X-Claim category which does not use any entity-type information to show the effectiveness of utilizing entity-type information for verifiable claim detection. Except for the X-Claim category, each word in the input sequence is identified with an entity type index $et_{i}$ according to one of the entity type categories. This information is fed into the entity type embedding layer to obtain a dense vector representation $ee_{i}$ for each word in the input sequence.
3.5 Entity-aware Claim Detection
Notation Description
$w_{i}$ $i^{th}$ word from the input sequence
$n$ Number of words in the input sequence
$we_{i}\in\mathbb{R}^{d_{w}}$ Word embedding of $i^{th}$ word
$d_{w}$ Dimension of word embedding
$et_{i}\in\mathbb{R}^{k}$ One-hot vector indicating entity type of $i^{th}$ word
$k$ Number of entity types
$ee_{j}\in\mathbb{R}^{d_{e}}$ Entity type embedding of $j^{th}$ entity type
$d_{e}$ Dimension of entity type embedding
$d_{p}$ Projection layer dimension
$P_{w}\in\mathbb{R}^{d_{w}\times d_{p}}$ Word embedding projection matrix
$P_{e}\in\mathbb{R}^{d_{e}\times d_{p}}$ Entity type embedding projection matrix
$h_{i}\in\mathbb{R}^{d_{p}}$ Projected input representation
$H$ Projected input sequence $(h_{1},h_{2},...h_{n})$
$se\in\mathbb{R}^{d_{p}}$ Sentence embedding
$C$ Number of class labels
$y\in\mathbb{R}^{C}$ Probability distribution predicted by the model for the input sequence
$\hat{y}\in\mathbb{R}^{C}$ Ground truth probability distribution of the input sequence
$EE\in\mathbb{R}^{k\times d_{e}}$ Entity embedding layer
$Attn$ Attention layer
Table 1: List of Key Notations
Given an input word sequence $(w_{1},w_{2},\ldots,w_{n})$, we extract the cross-lingual word embeddings $(we_{1},we_{2},\ldots,we_{n})$ , and the entity type index $(et_{1},et_{2},\ldots,et_{n})$. Here, $et_{i}\in\mathbb{R}^{k}$ represents the entity type of a word as a one-hot vector, where $k$ indicates the number of entity types. These entity-type indices are used to extract entity-type embeddings as follows,
$EE(et_{1},et_{2},...et_{n})=(ee_{1},ee_{2},...ee_{n})$ (1)
Once the entity-type embeddings are obtained, we project both word embeddings and entity-type embeddings into a common vector space and add the projected vectors at the word level as follows,
$h_{i}=P_{w}\cdot we_{i}+P_{e}\cdot ee_{i}\in\mathbb{R}^{d_{p}}$ (2)
This projection layer is responsible for converting the input features into task-specific, low-dimensional vectors and incorporating entity-type information into the word embedding.
To attend to the important words indicating the verifiability of claims, we apply scaled dot-product attention (Vaswani et al., 2017) to the projected input sequence $H=(h_{1},h_{2},...h_{n})$ as follows,
$Attn(H)=softmax(\frac{HH^{T}}{\sqrt{d_{p}}})\cdot H$ (3)
The attention layer $Attn$ generates attended context vectors for each element in the input sequence. The mean value of this context vector is obtained as the sentence representation $se$ as follows,
$se=\frac{1}{n}\sum Attn(H)\in\mathbb{R}^{2d_{p}}$ (4)
Finally, the sentence embedding $se$ is fed into a fully connected layer followed by a softmax function to generate a probability distribution $y$ over the output classes as follows,
$out=W_{o}se+b_{o}\in\mathbb{R}^{C}$ (5)
$y_{j}=\frac{exp(out_{j})}{\sum_{j}^{C}exp(out_{j})}\in\mathbb{R}^{C}$ (6)
Given the true label distribution $\hat{y}\in\mathbb{R}^{C}$, the entity type embedding layer, bidirectional GRU, and attention layer are trained by optimizing the cross-entropy loss.
4 Experiment Setting
4.1 Dataset
We use CheckThat! 2022 Nakov et al. (2022) dataset for training and testing the claim detection models. This dataset contains tweets related to COVID-19, written in 5 languages. Further, it contains four partitions: Train, Dev, Dev-test, and Test. We used the Train partition of all languages except Arabic for training and used the Dev partition as validation data. The Arabic data was excluded during the training as the initial experiments yielded poor performance in all languages including Arabic. Further, the Dev-Test partition was used to find the optimal parameters.
To evaluate the models’ ability to identify claims written in a broad range of unseen languages, we generated synthetic data from the CheckThat! Test partition for 18 unseen languages. For each, we randomly sampled 250 instances and translated them using Microsoft Azure AI translator Junczys-Dowmunt (2019), limiting samples to match the test dataset size. Translation quality was assessed on the FLORES-200 benchmark Team et al. (2022), with a CHRF score of 54.35. Further details are in Appendix B.2.
To evaluate the generalization capability of the proposed model in identifying claims discussing unseen topics as well as unseen language, we report the performance in Kazemi 2021 Kazemi et al. (2021). This dataset contains WhatsApp tipline and social group messages written in 5 languages. Refer to Appendix A.3 for the statistics of the datasets.
Model Accuracy Precision Recall F1-Score
CheckThat! 2022
mBERT 0.722 $\pm$ 0.006 0.723 $\pm$ 0.006 0.722 $\pm$ 0.006 0.715 $\pm$ 0.007
XLM-R 0.669 $\pm$ 0.007 0.695 $\pm$ 0.006 0.669 $\pm$ 0.007 0.658 $\pm$ 0.009
mT5 0.681 $\pm$ 0.004 0.685 $\pm$ 0.004 0.681 $\pm$ 0.004 0.679 $\pm$ 0.004
Llama2 0.776 $\pm$ 0.001 0.782 $\pm$ 0.001 0.776 $\pm$ 0.001 0.771 $\pm$ 0.001
Mistral 0.769 $\pm$ 0.003 0.784 $\pm$ 0.003 0.769 $\pm$ 0.003 0.761 $\pm$ 0.003
Phi3 0.744 $\pm$ 0.001 0.75 $\pm$ 0.001 0.744 $\pm$ 0.001 0.738 $\pm$ 0.001
X-Claim 0.747 $\pm$ 0.005 0.751 $\pm$ 0.005 0.747 $\pm$ 0.005 0.745 $\pm$ 0.005
EXN-Claim 0.754 $\pm$ 0.005 0.758 $\pm$ 0.006 0.754 $\pm$ 0.005 0.752 $\pm$ 0.006
EXP-Claim 0.755 $\pm$ 0.007 0.759 $\pm$ 0.007 0.755 $\pm$ 0.007 0.754 $\pm$ 0.007
Synthetic Data
mBERT 0.687 $\pm$ 0.009 0.704 $\pm$ 0.008 0.687 $\pm$ 0.009 0.679 $\pm$ 0.01
XLM-R 0.662 $\pm$ 0.008 0.68 $\pm$ 0.007 0.662 $\pm$ 0.008 0.656 $\pm$ 0.009
mT5 0.659 $\pm$ 0.007 0.662 $\pm$ 0.007 0.659 $\pm$ 0.007 0.658 $\pm$ 0.007
Llama2 0.724 $\pm$ 0.002 0.749 $\pm$ 0.002 0.724 $\pm$ 0.001 0.707 $\pm$ 0.002
Mistral 0.718 $\pm$ 0.002 0.759 $\pm$ 0.003 0.718 $\pm$ 0.002 0.701 $\pm$ 0.003
Phi3 0.691 $\pm$ 0.002 0.735 $\pm$ 0.005 0.691 $\pm$ 0.002 0.664 $\pm$ 0.002
X-Claim 0.719 $\pm$ 0.007 0.724 $\pm$ 0.006 0.719 $\pm$ 0.007 0.718 $\pm$ 0.007
EXN-Claim 0.73 $\pm$ 0.008 0.735 $\pm$ 0.007 0.73 $\pm$ 0.008 0.729 $\pm$ 0.008
EXP-Claim 0.731 $\pm$ 0.009 0.737 $\pm$ 0.008 0.731 $\pm$ 0.009 0.73 $\pm$ 0.009
Kazemi 2021 Data
mBERT 0.716 $\pm$ 0.007 0.725 $\pm$ 0.006 0.716 $\pm$ 0.007 0.702 $\pm$ 0.009
XLM-R 0.687 $\pm$ 0.011 0.752 $\pm$ 0.004 0.687 $\pm$ 0.011 0.686 $\pm$ 0.012
mT5 0.681 $\pm$ 0.004 0.725 $\pm$ 0.004 0.681 $\pm$ 0.004 0.678 $\pm$ 0.005
Llama2 0.633 $\pm$ 0.003 0.5 $\pm$ 0.001 0.633 $\pm$ 0.001 0.613 $\pm$ 0.004
Mistral 0.603 $\pm$ 0.004 0.747 $\pm$ 0.005 0.603 $\pm$ 0.004 0.55 $\pm$ 0.005
Phi3 0.546 $\pm$ 0.001 0.744 $\pm$ 0.001 0.546 $\pm$ 0.001 0.455 $\pm$ 0.001
X-Claim 0.788 $\pm$ 0.007 0.817 $\pm$ 0.005 0.788 $\pm$ 0.007 0.789 $\pm$ 0.007
EXN-Claim 0.793 $\pm$ 0.008 0.813 $\pm$ 0.006 0.793 $\pm$ 0.008 0.796 $\pm$ 0.008
EXP-Claim 0.794 $\pm$ 0.008 0.813 $\pm$ 0.006 0.794 $\pm$ 0.008 0.797 $\pm$ 0.007
All
mBERT 0.698 $\pm$ 0.008 0.711 $\pm$ 0.007 0.698 $\pm$ 0.008 0.689 $\pm$ 0.009
XLMR 0.668 $\pm$ 0.008 0.695 $\pm$ 0.006 0.668 $\pm$ 0.008 0.661 $\pm$ 0.01
mT5 0.667 $\pm$ 0.006 0.677 $\pm$ 0.006 0.667 $\pm$ 0.006 0.665 $\pm$ 0.006
Llama2 0.717 $\pm$ 0.002 0.755 $\pm$ 0.001 0.717 $\pm$ 0.002 0.702 $\pm$ 0.002
Mistral 0.707 $\pm$ 0.003 0.761 $\pm$ 0.004 0.707 $\pm$ 0.003 0.685 $\pm$ 0.003
Phi3 0.674 $\pm$ 0.002 0.74 $\pm$ 0.004 0.674 $\pm$ 0.002 0.64 $\pm$ 0.002
X-Claim 0.736 $\pm$ 0.006 0.746 $\pm$ 0.006 0.736 $\pm$ 0.006 0.735 $\pm$ 0.007
EXN-Claim 0.746 $\pm$ 0.007 0.753 $\pm$ 0.007 0.746 $\pm$ 0.007 0.745 $\pm$ 0.008
EXP-Claim 0.747 $\pm$ 0.008 0.754 $\pm$ 0.007 0.747 $\pm$ 0.008 0.746 $\pm$ 0.009
Table 2: Dataset-level Performance of the Claim Detection Models
4.2 Evaluation metrics
We report accuracy, precision, recall, and F1-score for ten fine-tuned claim detection models, presenting their averages and standard deviations. To account for varying ratios of verifiable and unverifiable claims across languages, we calculated weighted scores at the language level, and the overall performance is obtained by averaging the language-level performance of the models.
4.3 Baseline models
Due to the lack of cross-lingual claim detection models, we use strong cross-lingual language models as baselines: mBERT (Alam et al., 2021b; Uyangodage et al., 2021; Panda and Levitan, 2021; Zengin et al., 2021; Hasanain and Elsayed, 2022), XLM-RoBERTa (XLM-R) (Alam et al., 2021b, a), and mT5 (Du et al., 2022), all fine-tuned on the CheckThat! Train partition. These widely used models have proven competitive in multilingual claim detection. We also include open-source large language models from Li et al. (2024)—Llama2-7B, Mistral-7B, and Phi3—fine-tuned on the same data for text generation.
Refer to Appendix D for resources used for the training, hyperparameters, the prompt used for LLMs, and the LLM fine-tuning process.
5 Results
5.1 Dataset-level Performance
Table 2 summarizes the performance of the claim detection models. Llama2 performs strongly on the CheckThat! 2022 Test dataset, while EXP-Claim achieves the highest accuracy, recall, and F1-score on both the synthetic and Kazemi datasets. The performance of large language models declines on these two datasets, highlighting challenges in handling linguistically diverse data.
Notably, the EX-Claim models and their ablated variants outperform both transformer-based baselines and large language models in overall evaluation, underscoring the value of learning task-specific representations that focus on claim-relevant words. Moreover, their consistent performance across diverse datasets—spanning multiple platforms and topics—suggests that our hypothesis holds generally, regardless of the claim’s nature.
Incorporating NER information (EXN-Claim vs. X-Claim) consistently improves performance across all four metrics. While the overall gain is modest (approximately 1%), our language-level analysis (Section 5.2) shows that NER integration either maintains or enhances model performance across most of the languages, with certain languages benefiting significantly from entity information and linguistic knowledge transfer from the NER and EL tools. Interestingly, incorporating entity popularity (EXN-Claim vs. EXP-Claim) did not lead to further performance improvement—likely because most entities identified by the NER model were already recognized as popular, limiting the potential for additional gains.
5.2 Language-level Gain
(a) Transformer-based Baselines
(b) X-Claim
(c) Large Language Models (LLMs)
Figure 2: Performance Gain in EXN-Claim compared to Three groups of Models
We analyze language-level performance gains from incorporating entity information into claim detection. Since there is no performance gain in EXP-Claim compared to EXN-Claim, we analyze the performance of EXN-Claim with three groups of models as shown in Figure 2. Refer to Appendix B.1 for the language-level F1-score of the models.
Figure 2(a) illustrates the performance gain of the EXN-Claim model over transformer-based baselines. Most languages exhibit consistent positive gains, except English (Kazemi dataset) and Vietnamese, which show slight drops (0.95–1.2%). The most substantial improvements are seen in Asian languages like Malayalam, Tamil, Hindi, Bengali, Kannada, and Gujarati. These results underscore the effectiveness of incorporating entity information and dynamically attending to claim-relevant words, leading to significant performance gains across a diverse set of languages.
Figure 2(b) shows the performance gain of the EXN-Claim model over its ablation variant, X-Claim, which excludes entity information. Interestingly, incorporating entity information either retains or improves model performance across most languages. The only exceptions are German and Thai, which exhibit a slight performance drop of approximately 0.97% to 1.1%. Notably, certain languages—such as Kannada, Ukrainian, Bengali, Czech, Italian, and French—experience the highest gains, ranging from 1.95% to 4.6%. These improvements can likely be attributed to the incorporation of entity information and the indirect linguistic knowledge transfer from the NER tool.
Similarly, we compare the performance gain of the EXN-Claim model with large language models (LLMs) in Figure 2(c). While LLMs outperform EXN-Claim in several high-resource languages, they perform notably poorly in certain low-resource languages—particularly Asian regional languages such as Malayalam, Kannada, Tamil, Gujarati, Bengali, and Hindi. The largest gains are in Malayalam (up to 70%) and Tamil (up to 60%), which include non-synthetic samples from the Kazemi dataset. Despite being smaller and more computationally efficient, EXN-Claim outperforms all three LLMs in 7 languages and surpasses at least one LLM in 16 out of the 27 languages evaluated.
Furthermore, we observe inconsistent gains across LLMs: Llama excels in Dutch, Phi3 in Hungarian, while Mistral underperforms in both. Similarly, Llama and Phi3 perform well in Korean, but Mistral struggles. These observations suggest that LLMs tend to excel only in specific languages, largely influenced by the composition of their training data. In contrast, these findings highlight EXN-Claim as a highly efficient model that consistently delivers superior performance across many low-resource languages, while remaining competitive with LLMs in high-resource scenarios, offering the best balance between overall performance gains and computational efficiency. Refer to Appendix D.3 for comparison of claim detection models based on the number of parameters, multilingual support, and training time.
6 Analysis
6.1 Cross-lingual Knowledge Transfer
The synthetic data comprises claims from the CheckThat! Test partition, written in 18 different languages. We use this dataset to assess the cross-lingual knowledge-transferring capabilities of the claim detection models. We compute the fraction of data instances in the Synthetic data which gets the same class prediction as its original data instance from the CheckThat! Test partition by a model. We categorized these results into correct and incorrect predictions, allowing us to evaluate the model’s ability to transfer knowledge across correct and wrong predictions.
Model Correct Prediction Wrong Prediction
mBERT 82.75% 64.83%
XLMR 81.72% 66.27%
mT5 79.12% 60%
Llama2 85.4% 70.1%
Mistral 86.87% 71.3%
Phi3 82.46% 65.5%
X-Claim 87.21% 71.73%
EXN-Claim 87.82% 68.48%
EXP-Claim 87.09% 66.51%
Table 3: Cross-lingual Knowledge Transfer Rate
Table 3 lists the knowledge transfer rate of claim detection models. Overall, the transformer-based baseline models exhibit the lowest transfer rate in correct and wrong predictions. In contrast, EX-Claim and its ablated variant achieve the highest transfer rate—reaching nearly 87%—for correct predictions. This highlights the effectiveness of learning context-aware claim representations with attention to claim-relevant words in promoting cross-lingual generalization. Furthermore, the inclusion of entity information through NER and EL (EXN-Claim and EXP-Claim) reduces the transfer of incorrect predictions by 3% to 5%. This indicates that entity-level knowledge and linguistic signals from NER and EL help prevent consistent misclassification of the same instances across different languages, thus improving model robustness.
6.2 Attending to Important Words
The attention layer enables the claim detection models to attend to important words in the input sequence before generating the sentence representation. We quantify this behavior of the models as the entropy of attention weights computed by the attention layer. We compute the average entropy of the test data by computing the mean value of entropy of all the input sequences.
Model Entropy
X-Claim 4.09
EX-Claim 3.519
EXP-Claim 3.322
Table 4: Entropy of Attention Layer
Table 4 lists the attention layer’s entropy of the claim detection models. We can see that the models utilizing entity information always yield lower entropy, especially when utilizing popularity, resulting in the lowest entropy among all. This shows that these models are concentrated on a few keywords in the input sequence to determine the verifiability of a claim. [^2] Refer to Appendix C.2 for attention weights generated for sample sentences.
6.3 Error Analysis
False Positive
S4 Months after receiving the second dose of the Pfizer COVID vaccine, I continue to suffer side effects including perfect health and having to deal with morons.
S5 How true is Congress party’s claim regarding shortfall in Government earnings
False Negative
S6 These are not photos of a man with an artificial heart
S7 Retirement Home NOT Raided By The FBI For Running Elderly Fight Club, NO 7 Arrests
S8 Gruesome Video Of A Man Hacked To Death In Bihar Falsely Shared As TMC ’Goons’ Attacking A BJP Worker
S9 Video Showing Suicide Bid Falsely Linked To CAA
Table 5: Misclassified Sample Sentences
We analyze the false positive and false negative predictions made by EX-Claim models. Table 5 lists some of the misclassified examples. We noticed that the appearance of disease names ($S4$), lengthy claims, nonrecognition of personal experience from pronouns’ presence ($S4$), and ambiguous fact expressions could cause false positive scenarios ($S5$).
We observed that misclassification of popular events expressing personal experience tends to be a common cause of false negatives. Refer to Sentences S6-S9 in Table 5 for sample false negative instances where the reach of the event makes it verifiable, yet the personal experience framing leads to misclassification. Further, some of these claims refer to other modalities such as images and videos in the expression of facts, making it further difficult when relying solely on textual information  [^3] Refer to Appendix C.4 showing the correlation between the confusion matrix and some of these features..
7 Conclusion
We present EX-Claim, the first cross-lingual claim detection model that integrates entity information for improved performance in detecting claims across multiple languages. Extensive experiments on three datasets comprising 27 languages reveal that EX-Claim stands out as the effective model with consistent performance gain across multiple languages compared to the baselines. Language-level performance analysis reveals that certain languages—particularly Asian regional languages—benefit significantly from the incorporation of entity information. This improvement is likely driven by the emphasis on key terms essential for claim detection and by the indirect transfer of linguistic knowledge from named entity recognition and entity linking tools. Further analysis of cross-lingual transfer rates and attention layer entropy supports this finding. Notably, EX-Claim achieves the highest cross-lingual knowledge transfer rate across correct predictions while minimizing the wrong predictions across claims written in the same language, demonstrating its robustness in cross-lingual generalization and transfer efficiency. Future work will focus on generalizing entity-type embeddings beyond training contexts and address the challenges contributing to false-positive and false-negative outcomes identified in our error analysis.
8 Limitations
The limitations of this work are as follows:
  • Training data limited to a single topic - Our training data consists solely of COVID-19 discussions. Although our model generalizes well across different datasets, including those focused on political topics, this limitation may still restrict the model’s broader generalization capabilities.
  • Dependence on NER and entity linking tools: The proposed solution relies on existing cross-lingual NER and entity linking tools to extract entity-type information. The performance of these tools directly impacts the overall effectiveness of our model, particularly in languages where NER or entity-linking tools may perform poorly or exceptionally well.
  • Training data limited to a single platform: Our training data is exclusively composed of tweets, which may cause the model to learn characteristics specific to this platform. While we demonstrate that our model can handle data from other platforms, such as WhatsApp tiplines in the Kazemi dataset, this limitation could still hinder the model’s ability to generalize across different types of social media content.
  • Exclusion of popular events not identified by a name: Since the system identifies and tracks the popularity of named entities, popular events not linked to any names are excluded. As previously discussed, this is a significant cause of false negative predictions.
Acknowledgments
This project is funded by the European Union and UK Research and Innovation under Grant No. 101073351 as part of Marie Skłodowska-Curie Actions (MSCA Hybrid Intelligence to monitor, promote, and analyze transformations in good democracy practices).
Appendix A Statistics
Figure 3: Average Number of Named Entities Identified in the Train Partition of CheckThat! 2022 Data
A.1 Statistics of Named Entities Recognized
MultiNERD identifies the following 15 types of fine-grained Named Entities.
  • Person (PER)
  • Organization (ORG)
  • Location (LOC)
  • Animal (ANIM)
  • Biological entity (BIO)
  • Celestial Body (CEL)
  • Disease (DIS)
  • Event (EVE)
  • Food (FOOD)
  • Instrument (INST)
  • Media (MEDIA)
  • Plant (PLANT)
  • Mythological entity (MYTH)
  • Time (TIME)
  • Vehicle (VEHI)
Figures 3 show the average number of named entities recognized in the Train partition of CheckThat! 2022 data. Notably, location entities appear more frequently across all languages. This could be due to the COVID-19 discussion in the training data referring to the status of the disease from different countries and locations. Additionally, the occurrences of disease, media, biological entities, and food are relatively higher compared to other rare entity types, likely due to the COVID-19-related discussions present in the dataset.
A.2 Statistics of Entities Linked
Figure 4: Number of Entities Linked with the Variation in EL Threshold in Train Partition of CheckThat! Data
Figure 5: Popular vs Unpopular Fine-grained Entities
Figure 6: Fraction of Popular and Unpopular Entities in Train Partition of CheckThat! Data
Figure 4 shows the number of named entities linked to a Wikipedia page with the increase in threshold. It can be observed that the reduction in the number of entities linked to a Wikipedia page with the increase in log probability threshold is not very significant until the threshold is increased to -0.15 (probability score of 0.861). This indicates that most of the NER identified are linked to a Wikipedia page with a very high probability.
Figure 6 shows the fraction of popular and unpopular entities determined with the log probability threshold of -0.15 (probability of 0.861). Notably, Arabic language data contains more fraction of entities identified as popular followed by English, and Dutch and Turkish have the lowest fraction among all. Further, the chances of fine-grained entities being popular range from 0.39 to 0.63, highlighting that nearly more than 40% of the entities recognized are linked as popular ones with a very high probability.
A.3 Statistics of the Datasets Used
A.3.1 CheckThat! Dataset
Arabic Bulgarian Dutch English Turkish Total
Train Verifiable 2,513 1,871 929 2,122 1,589 14,032
Unverifiable 1,118 839 1,021 1,202 828
Total 3,631 2,710 1,950 3,324 2,417
Dev Verifiable 235 177 72 195 150 1,300
Unverifiable 104 74 109 112 72
Total 339 251 181 307 222
Dev-Test Verifiable 691 519 252 574 438 3,837
Unverifiable 305 217 282 337 222
Total 996 736 534 911 660
Test Verifiable 682 199 608 149 303 3,698
Unverifiable 566 130 750 102 209
Total 1,248 329 1,358 251 512
Table 6: Statistics of the CheckThat! 2022 Verifiable Claim Detection Dataset
Table 6 presents the statistics of the CheckThat! 2022 dataset. Compared to other languages, the number of data instances is low in Dutch across all partitions except Test. Further, the ratio between verifiable and unverifiable claims is equal for Dutch languages, whereas other languages contain more verifiable statements than unverifiable (roughly 2:1 ratio).
A.3.2 Kazemi Dataset
Bengali English Hindi Malayalam Tamil Total
Verifiable 329 535 407 707 197 4,017
Unverifiable 593 388 465 263 133
Total 922 923 872 970 330
Table 7: Statistics of the Kazemi 2021 Verifiable Claim Detection Dataset
Table 7 presents the statistics of the Kazemi dataset. The dataset contains WhatsApp tiplines and social media posts about topics such as politics and COVID-19.
Appendix B Results
Baseline Entity Type Categories
Language
mBERT
XLM-R
mT5
Llama2
Mistral
Phi3
X-Claim
EXN-Claim
EXP-Claim
Arabic 0.692 0.654 0.677 0.752 0.736 0.718 0.731 0.733 0.735
Bengali 0.725 0.655 0.556 0.697 0.642 0.552 0.713 0.748 0.764
Bulgarian 0.783 0.633 0.729 0.835 0.815 0.812 0.782 0.79 0.796
Czech 0.707 0.647 0.72 0.808 0.786 0.755 0.741 0.774 0.769
Dutch 0.68 0.698 0.634 0.72 0.657 0.663 0.716 0.717 0.718
English (Kaz) 0.738 0.769 0.748 0.764 0.75 0.719 0.755 0.757 0.755
English (CT) 0.724 0.658 0.688 0.769 0.777 0.755 0.75 0.762 0.761
French 0.638 0.698 0.658 0.736 0.739 0.738 0.703 0.722 0.725
German 0.751 0.714 0.722 0.808 0.822 0.816 0.789 0.778 0.775
Greek 0.677 0.64 0.71 0.697 0.687 0.679 0.735 0.747 0.759
Gujarati 0.661 0.669 0.649 0.457 0.582 0.455 0.77 0.77 0.751
Hindi 0.631 0.695 0.613 0.792 0.809 0.6 0.766 0.767 0.769
Hungarian 0.667 0.651 0.665 0.757 0.727 0.637 0.68 0.683 0.682
Italian 0.721 0.647 0.704 0.764 0.745 0.736 0.709 0.731 0.723
Japanese 0.716 0.658 0.693 0.759 0.771 0.74 0.715 0.732 0.746
Kannada 0.535 0.567 0.579 0.371 0.459 0.342 0.628 0.674 0.676
Korean 0.691 0.635 0.623 0.69 0.72 0.653 0.712 0.718 0.709
Marathi 0.588 0.64 0.603 0.643 0.532 0.46 0.668 0.666 0.664
Malayalam 0.79 0.564 0.685 0.374 0.151 0.141 0.845 0.84 0.835
Persian 0.673 0.669 0.622 0.713 0.667 0.602 0.704 0.716 0.718
Russian 0.709 0.667 0.641 0.766 0.764 0.734 0.715 0.724 0.724
Spanish 0.738 0.693 0.683 0.779 0.735 0.77 0.787 0.783 0.784
Swedish 0.707 0.688 0.654 0.772 0.779 0.76 0.777 0.778 0.787
Tamil 0.623 0.744 0.789 0.439 0.398 0.264 0.864 0.866 0.861
Thai 0.628 0.648 0.644 0.716 0.698 0.673 0.726 0.716 0.715
Turkish 0.699 0.648 0.664 0.775 0.82 0.745 0.748 0.759 0.758
Ukrainian 0.688 0.597 0.604 0.764 0.724 0.719 0.644 0.688 0.707
Vietnamese 0.727 0.672 0.676 0.735 0.694 0.689 0.715 0.717 0.729
Table 8: F1-Score of the Claim Detection Models at Language-level
B.1 Language-level Performance
Table 8 lists the language-level F1-score of the claim detection models across 27 languages present in the test datasets.
B.2 Machine Translation Performance
Since the synthetic dataset was generated using machine translation, we assess the quality of the translation model—Microsoft Azure AI Translator Junczys-Dowmunt (2019). For this evaluation, we utilize FLORES-200 Team et al. (2022), a multilingual benchmark dataset that supports English-to-200+ language translation. To approximate the conditions of our synthetic dataset, we sampled sentence pairs from the five source languages represented in the CheckThat! training set to 18 target languages present in the synthetic dataset. For each target language, we selected 250 translation pairs, resulting in a total of 4,500 sentence pairs. This sampling strategy is designed to mimic the structure of our synthetic dataset, where each target language includes 250 translated claims.
Metric Score
BLEU 26.23
CHRF 54.35
CHRF++ 51.63
Table 9: Machine Translation Performance of Microsoft Azure AI translator
Table 9 presents the performance of the machine translation model on the sampled translation pairs, evaluated using multiple translation quality metrics. The model achieves competitive scores on the character-based metrics CHRF and CHRF++, with the CHRF++ score notably surpassing the baseline results reported in the original FLORES-200 dataset paper Team et al. (2022).
Appendix C Analysis
C.1 Entity-type Embedding
Figure 7: Two-dimensional embedding visualization of EXN-Claim Model
Unlike the cross-lingual word embedding which is directly obtained from the pre-trained model, EX-Claim learns a dense vector representation for each entity type. Figure 7 shows the 2D visualization of the entity type embeddings learned by the EXN-Claim model. The 2D vector representation of the embeddings is obtained using the t-SNE [^4] https://scikit-learn.org/stable/modules/generated/sklearn.manifold.TSNE.html dimension reduction technique. Further, the embedding representations were clustered using the KMeans clustering technique (Sinaga and Yang, 2020), and the number of optimal clusters was chosen using the Silhouette score (Rousseeuw, 1987).
The EXN-Claim model learns various groups of entity types. Notably, this model clusters Animal, Disease, Person, Biological Entity, Disease, Plan and Mythological Entity into a single group, likely influenced by the COVID-19 discussions present in the training data. This observation highlights the model’s tendency to learn context-specific entity-type embeddings during training.
C.2 Attention Weights for Sample Sentences
Figure 9 shows the heatmaps of the attention weights computed by different claim detection models for the following three sentences.
  • S10 - Old video of cop beating woman in Gawalior shared as Delhi police brutality
  • S11 - Donald Trump fund raising email takes CNN anchor’s comments out of context
  • S12 - Trump withdraws US from the Iran nuclear deal
It can be observed that EX-Claim attends to important words in the sentences, whereas X-Claim generates attention weights distributed across most of the words. Further, an interesting pattern of named entities attending to themselves or other named entities more compared to other words can be observed across all three sentences.
C.3 Ablation Study
We conduct an ablation study to identify the impact of the key components of EX-Claim with the following variations.
  • No projection layer (NPL) - The projection layer is omitted, and attention is applied directly to the embedding representations.
  • No attention layer (NAL) - The attention layer is removed, hence the projected vector representations are aggregated into a sentence embedding by computing their mean.
  • No embedding layer (NEL) - The embedding layer is removed, and the entity information is represented as one-hot vectors.
Model F1-Score
X-Claim 0.735
X-Claim - NPL 0.688
X-Claim - NAL 0.714
Table 10: Ablation Results of X-Claim Model
The NLP and NAL variations are applied to X-Claim, and Table 10 presents the overall F1-score of these models. The results demonstrate that the projection layer is crucial for learning task-specific input representations, as its removal leads to a 4.7% drop in overall performance. Likewise, the attention layer enhances the model’s performance, contributing to a 2.1% improvement.
Model F1-Score
EXN-Claim 0.745
EXN-Claim - NEL 0.733
EXP-Claim 0.746
EXP-Claim- - NEL 0.733
Table 11: Ablation Results of EX-Claim Models
Table 11 lists the performance of EX-Claim models with NEL variation. It can be seen, that the performance declines by approximately 1.1% - 1.3% when entity information is represented as one-hot vectors, suggesting that the models struggle to effectively utilize entity information when it is provided as a one-hot vector representation.
C.4 Correlation between Confusion Matrix and Selected Features
Figure 8: Correlation of Sample Features vs Confusion Matrix
Figure 8 presents the correlation scores between the true positive, true negative, false positive, and false negative values and selected features. It can be observed that the length of the claim is positively correlated with true positive and false positive rates, showing the models tend to assume lengthy claims as verifiable.
Similarly, the number of disease names appearing in a claim is positively correlated with a false positive rate. This could be the COVID-19 discussion in the training data making the model to classify the claims with disease names as verifiable. Further, the number of media names is positively correlated with both true negative and false negative rates as discussed in Section 6.3. This suggests that references to media, such as images and videos, can either enhance or hinder verifiability, depending on the context and reach.
(a) S4 : EXN-Claim
(b) S4 : X-Claim
(c) S11 : EXN-Claim
(d) S11 : X-Claim
(e) S12 : EXP-Claim
(f) S12 : X-Claim
Figure 9: Heatmap of Attention Weights Generated by Claim Detection Models for sentences S10, S11, and S12
Appendix D Implementation
Model # Parameters Multilingual? Training time
Transformer-based Baselines mBERT 110M ✓ Yes (104 langs) 20 sec
XLM-R 270M ✓ Yes (100 langs) 20 sec
mT5 Large Encoder 600M ✓ Yes (101 langs) 20 sec
LLMs LLaMA 2 7B Limited 3 hours & 16 mins
Mistral 7B Limited 3 hours & 25 mins
Phi-3 3.8B Limited 2 hours & 12 mins
Entity-aware Models X-Claim 197K ✓ Yes (100 langs) 20 sec
EXN-Claim 201K
✓ Yes (100 langs
+ NER model langs)
20 sec
EXP-Claim 206K
✓ Yes (100 langs
+ NER model langs
& EL model langs)
21 sec
Table 12: Comparison of Model Parameters, Multilingual Support, and Training Time
D.1 Resources
All the experiments were conducted using Queen Mary’s Apocrita HPC facility, supported by QMUL Research-IT King et al. (2017). Specifically, 1 GPU (Volta V100 or Ampere A100) with 8 CPU cores, each composed of 11 GB memory was used to train and test all the models.
D.2 Models
We fine-tuned the mBERT [^5] https://huggingface.co/google-bert/bert-base-multilingual-cased, XLM-R [^6] https://huggingface.co/FacebookAI/xlm-roberta-base, and mT5 models published in HuggingFace to obtain the transformer-based baseline models. Especially, mT5 is available in various sizes. We used mT5-large [^7] https://huggingface.co/google/mt5-large as this model was the largest model among the variations which we could accommodate with the same resources utilized for training and testing other models. Similarly, the open-source large language models (LLMs) available in HuggingFace were used to obtain the LLM baselines, Llama2888https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, Mistral 7B [^9] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2, and Phi3101010https://huggingface.co/microsoft/Phi-3-mini-4k-instruct. We used the same experimental setting of Li et al. (2024). The prompt 1 used to fine-tune LLMs is a modified version of Li et al. (2024).
### Instruction:
Evaluate whether the input text contains
information or claims that can be verified
through fact-checking. If the statement presents
assertions, facts, or claims, respond with ’Yes’.
If the statement is purely opinion-based, trivial,
or does not contain any verifiable information or
claims, respond with ’No’.
### Input Sentence: <input sentence>
### Response: <Yes/No>
Prompt 1 Verifiable Claim Detection
All the models were fine-tuned by extracting the vector representation of the special token from the final layer and applying a fully connected layer on it for classification. We used only the encoder of the mT5 model for obtaining vector representation of input words. For Named Entity Recognition (NER), we used the resources published for MultiNERD [^11] https://github.com/Babelscape/multinerd/tree/master for fine-tuning the fine-grained NER model. Similarly, the mGENRE [^12] https://huggingface.co/facebook/mgenre-wiki model published in HuggingFace was used for Entity Linking (EL).
D.3 Model Comparison
Table 12 compares claim detection models based on the number of parameters, multilingual support, and training time. Both EXN-Claim and EXP-Claim incur an additional average latency of 5.2 ms per sentence for named entity recognition. Furthermore, EXP-Claim requires an average of 0.15 seconds per entity to perform entity linking.
D.4 Hyperparameters
Hyperparameter Value
Projection Size 256
Learning Rate 3e-5
Entity Embedding Size 128 (X, EX), 256 (EXP)
Entity Linking Threshold -0.15
Table 13: Hyperparameters Used
Table 13 lists the hyperparameters used. Following Devlin et al. (2019), we set the maximum length of the input sequence to 128, and use Adam as the optimizer. We train the models for 30 epochs and choose the model with the lowest validation loss as the best model.